# PUP 3D-GS: Principled Uncertainty Pruning for 3D Gaussian Splatting

### v1

@@ Abstract
Recent advancements in novel view synthesis have enabled real-time rendering speeds and high reconstruction accuracy. 3D Gaussian Splatting (3D-GS), a foundational point-based parametric 3D scene representation, models scenes as large sets of 3D Gaussians. Complex scenes can comprise of millions of Gaussians, amounting to large storage and memory requirements that limit the viability of 3D-GS on devices with limited resources. Current techniques for compressing these pretrained models by pruning Gaussians rely on combining heuristics to determine which ones to remove. In this paper, we propose a principled spatial sensitivity pruning score that outperforms these approaches. It is computed as a second-order approximation of the reconstruction error on the training views with respect to the spatial parameters of each Gaussian. Additionally, we propose a multi-round prune-refine pipeline that can be applied to any pretrained 3D-GS model without changing the training pipeline. After pruning $88.44\%$ of the Gaussians, we observe that our PUP 3D-GS pipeline increases the average rendering speed of 3D-GS by 2.65×2.65$\times$2.65 × while retaining more salient foreground information and achieving higher image quality metrics than previous pruning techniques on scenes from the Mip-NeRF 360, Tanks & Temples, and Deep Blending datasets.
Figure 1: We prune the 3D-GS reconstruction of the Deep Blending playroom scene from $2.6$M Gaussians to $0.3$M Gaussians using our PUP 3D-GS pipeline and LightGaussian – a recent, high-performing post-hoc pruning pipeline for pretrained 3D-GS models. PUP 3D-GS retains more fine details while increasing the rendering speed from $121.81$ FPS to $273.29$ FPS, a $\boldsymbol{2.24\times}$ increase over the base 3D-GS model. In comparison, LightGaussian loses substantially more fine details and reaches a slower rendering speed of $233.44$ FPS. A detailed comparison of PUP 3D-GS and LightGaussian can be found in Section 5.3.2; additional per-scene metrics are recorded in Appendix A.1.
@@ 1 Introduction
Novel view synthesis aims to render views from new viewpoints given a set of 2D images. Neural Radiance Fields (NeRFs) [15] use volume rendering to represent the 3D scene using a multilayer perceptron that can be used to render novel views. Although NeRF and its variants achieve high-quality reconstructions, they have slow inference time and require several seconds to render a single image. 3D Gaussian Splatting (3D-GS) [9] has recently emerged as a faster alternative to NeRF, achieving real-time rendering on modern GPUs while producing comparable image quality. It represents scenes using a large set of 3D Gaussians with independent location, shape, and appearance parameters. Since they typically consist of millions of Gaussians, 3D-GS scenes often have high storage and memory requirements that limit their viability on devices with limited resources.
Several recent works propose techniques that reduce the storage requirements of 3D-GS models, including heuristics for pruning Gaussians that have an insignificant contribution to the rendered images [3, 4, 12, 18, 11]. In this work, we propose a more mathematically principled approach for deciding which Gaussians to prune from a 3D-GS scene. We introduce a computationally feasible sensitivity score that is derived from the Hessian of the reconstructed error on the training images in a converged scene. Following the methodology in LightGaussian [3], we prune the scene using our sensitivity score and then fine-tune the remaining Gaussians. Our experiments demonstrate that this process can remove up to $66\%$ of the Gaussians in the base 3D-GS scene while maintaining nearly identical image quality metrics.
Moreover, our approach enables a second consecutive round of pruning and fine-tuning. As shown in Figure 1, our multi-round pruning approach can remove over $88\%$ of the Gaussians in the base 3D-GS scene while surpassing previous heuristics-based methods like LightGaussian in both rendering speed and image quality. Our work is orthogonal to several other methods that modify the training framework of 3D-GS or apply quantization-based techniques to reduce its disk storage. We show that our PUP 3D-GS pipeline can be used in conjunction with these techniques to further improve performance and compress the model.
@@
    A post-hoc pruning pipeline, PUP 3D-GS, that can be applied to any pretrained 3D-GS model without changing its training pipeline.
@@
    A multi-round prune-refine approach that produces higher fidelity reconstructions than equivalent single-round prune-refining.
@@ 2.1 Uncertainty Estimation
Several works estimate uncertainty in NeRF-based approaches that arise from sources like transient objects, camera model differences, and lightning changes [8, 19, 20]. Other works focus on uncertainty caused by occlusion or sparsity of training views [14]. These approaches rely on an ensemble of models [24] or variational inference and KL Divergence [23, 22], requiring intricate changes to the training pipeline to model uncertainty.
BayesRays [5] uses Fisher information for post-hoc uncertainty quantification in NeRF-based approaches. Since NeRFs represent the scene as a continuous 3D function, they compute its Hessian over a hypothetical perturbation field to estimate uncertainty. In contrast, our approach focuses on 3D Gaussian Splats [9] instead of NeRF [14], and we directly compute the Fisher information using gradient information without relying on a hypothetical perturbation field.
@@ 2.2 Pruning Gaussian Splat Models
A recent pruning-based method, Compact-3DGS [11], proposes a learnable masking strategy to prune small, transparent Gaussians during training. EAGLES [4], which we apply our pipeline to in Appendix A.8, prunes Gaussians based on the least total transmittance per Gaussian. They also begin with low-resolution images, progressively increasing image resolution to reduce Gaussian densification during training, then quantize several attributes to reduce disk storage. LightGaussian [3] computes a global significance score for each Gaussian with heuristics, uses that score to prune the least significant Gaussians, and finally uses quantization to further reduce storage requirements.
3 Background
3.1 3D Gaussian Splatting
3D Gaussian Splatting (3D-GS) is a point-based novel view synthesis technique that uses 3D Gaussians to model the scene. The attributes of the Gaussians are optimized over input training views given by a set of camera poses $\mathcal{P}_{gt}=\{\phi_{i}\in{}^{3\times 4}\}_{i=1}^{K}$ and corresponding ground truth training images $\mathcal{I}_{gt}=\{\boldsymbol{I}_{i}\in{}^{H\times W}\}_{i=1}^{K}$. A sparse point cloud of the scene generated by Structure from Motion (SfM) over the training views is used to initialize the Gaussians. At fixed intervals during training, a Gaussian densification step is applied to increase the number of Gaussians in areas of the model where small-scale geometry is insufficiently reconstructed.
Each 3D Gaussian $\mathcal{G}_{i}$ is independently parameterized by position $x_{i}\in{}^{3}$, scaling $s_{i}\in{}^{3}$, rotation $r_{i}\in{}^{4}$, base color $c_{i}\in{}^{3}$, view-dependent spherical harmonics $h_{i}\in{}^{15\times 3}$, and opacity $\alpha_{i}\in\real$. From these, we define the set of all Gaussian parameters as:
@@ 3 Background
For a given camera pose $\phi$, a differential rasterizer renders 2D image $I_{\mathcal{G}}(\phi)$ by projecting all Gaussians observed from $\phi$ onto the image plane. The color of each pixel in $I_{\mathcal{G}}(\phi)$ is given by the blending of the $\mathcal{N}$ ordered Gaussians that overlap it:
$C=\sum_{i\in\mathcal{N}}\tilde{c}_{i}\tilde{\alpha}_{i}\prod_{j=1}^{j-1}(1-\tilde{\alpha}_{j}),$ (3)
where $\tilde{c}_{i}$ represents the view-dependent color calculated from the camera pose $\phi$ and the optimizable per-Gaussian color $c_{i}$ and spherical harmonics $h_{i}$, and $\tilde{\alpha_{i}}$ and $\tilde{\alpha_{j}}$ represent the projected opacities.
@@
where the first term is an L1 residual loss and the second term is an SSIM loss.
@@ 4 Method
We define uncertainty in 3D-GS as the amount that a Gaussian’s parameters, such as its location and scale, can be perturbed without affecting the reconstruction loss over the input views. Concretely, this is the sensitivity of the error over the input views to that particular Gaussian. Given a loss function $L:{}^{\mathcal{G}}\to\real$ that takes the set of Gaussians $\mathcal{G}$ as inputs and outputs an error value over the set of training views, this sensitivity can be captured by the Hessian $\nabla^{2}_{\mathcal{G}}L$.
In the following subsections, we demonstrate how to compute this sensitivity and use it to decide which Gaussians to prune from the model. Directly computing the full Hessian matrix is intractable due to memory constraints. We demonstrate how to obtain a close estimate of the Hessian via a Fisher approximation in Section 4.1. We also find that only a block-wise approximation of the Hessian parameters is needed to quantify Gaussian sensitivity in Section 4.2, and that computing this over image patches instead of pixels reduces compute and is sufficient for pruning in Section 4.3. Finally, we find that multiple rounds of pruning and fine-tuning improves our performance over one-shot pruning and fine-tuning, giving our full PUP 3D-GS pipeline in Section 4.4.
4.1 Fisher Approximation
@@
On a converged 3D-GS model, the $||I_{\mathcal{G}}-I_{gt}||_{1}$ residual term of Equation 4 approaches zero, causing the second order term $(I_{\mathcal{G}}(\phi)-I_{gt})\nabla_{\mathcal{G}}^{2}I_{\mathcal{G}}$ in Equation 6 to vanish. This leaves us with:
$\nabla_{\mathcal{G}}^{2}L_{2}=\nabla_{\mathcal{G}}I_{\mathcal{G}}\nabla_{\mathcal{G}}I_{\mathcal{G}}^{T},$ (7)
which is known as a Fisher approximation of the Hessian [5, 7].
Note that $\nabla_{\mathcal{G}}I_{\mathcal{G}}$ is the gradient over only the reconstructed images. This means that our approximation only depends on the input poses $\mathcal{P}_{gt}$ and not the input images $\mathcal{I}_{gt}$. However, our Fisher approximation of $\nabla_{\mathcal{G}}^{2}L_{2}$ is still computationally expensive as it requires a sum over all per-pixel Fisher approximations of the scene. We describe how we make this more efficient in Section 4.3.
@@ 4.2 Sensitivity Pruning Score
The full Hessian over the model parameters $\nabla_{\mathcal{G}}^{2}L\in{}^{\mathcal{G}\times\mathcal{G}}$ is quadratically large. However, we find that using only a fraction of these parameters is sufficient for an effective sensitivity pruning score.
@@
$\boldsymbol{H}_{i}=\nabla_{\mathcal{G}_{i}}I_{\mathcal{G}_{i}}\nabla_{\mathcal{G}_{i}}I_{\mathcal{G}_{i}}^{T}.$ (8)
Intuitively, each of these block Hessians measures the impact on the reconstruction error when we perturb the parameters of only the Gaussian it represents. To turn each Hessian $\boldsymbol{H}_{i}$ into a single value sensitivity score ${\tilde{U}}_{i}\in\real$, we take the log determinant of $\boldsymbol{H}_{i}$:
${\tilde{U}}_{i}=\log|\boldsymbol{H}_{i}|=\log|\nabla_{\mathcal{G}_{i}}I_{\mathcal{G}_{i}}\nabla_{\mathcal{G}_{i}}I_{\mathcal{G}_{i}}^{T}|.$ (9)
This captures the relative impact over all of the parameters of Gaussian $\mathcal{G}_{i}$ on the reconstruction error. Specifically, it is a relative volume measure of the basin around the Gaussian parameters $\mathcal{G}_{i}$ in the second-order approximation of the function describing their impact on the reconstruction error.
We find that we can restrict the per Gaussian parameters even further and consider only the spatial mean $x_{i}$ and scaling $s_{i}$ parameters for an effective sensitivity score. Our final sensitivity pruning score $U_{i}\in\real$ is:
$U_{i}=\log|\nabla_{x_{i},s_{i}}I_{\mathcal{G}_{i}}\nabla_{x_{i},s_{i}}I_{\mathcal{G}_{i}}^{T}|.$ (10)
We speculate that only the spatial mean and scaling parameters are needed because they capture the projective geometric invariances that exist between the 3D scene and the input views as shown in Figure 2. We ablate the choice of using spatial parameters in Section 6.2.
Figure 3: Spearman correlations between the $1\times 1$ pixel-wise sensitivity score and the $4\times 4$, $8\times 8$, and $64\times 64$ patch-wise sensitivity scores across all Gaussians in the Tanks & Temples truck scene. Notice that the correlation remains very high even with a large patch size of $64\times 64$.
Table 1: Mean PSNR, SSIM, LPIPS, FPS, and point cloud size for the Mip-NeRF 360 dataset at each stage of our PUP 3D-GS pipeline. The operation in each row is applied cumulatively to all of the following rows. Per-scene metrics across all datasets are recorded in Appendix A.1.
[…]
3D-GS 27.47 0.8123 0.2216 95.59 746.46
+ Prune $66\%$ 25.60 0.7836 0.2497 178.23 253.8
+ Refine 27.50 0.8122 0.2281 153.27 253.8
+ Prune $66\%$ 20.35 0.6718 0.3441 310.73 86.29
+ Refine 26.83 0.792 0.2688 244.07 86.29

### v3

@@ Abstract
Recent advances in novel view synthesis have enabled real-time rendering speeds with high reconstruction accuracy. 3D Gaussian Splatting (3D-GS), a foundational point-based parametric 3D scene representation, models scenes as large sets of 3D Gaussians. However, complex scenes can consist of millions of Gaussians, resulting in high storage and memory requirements that limit the viability of 3D-GS on devices with limited resources. Current techniques for compressing these pretrained models by pruning Gaussians rely on combining heuristics to determine which Gaussians to remove. At high compression ratios, these pruned scenes suffer from heavy degradation of visual fidelity and loss of foreground details. In this paper, we propose a principled sensitivity pruning score that preserves visual fidelity and foreground details at significantly higher compression ratios than existing approaches. It is computed as a second-order approximation of the reconstruction error on the training views with respect to the spatial parameters of each Gaussian. Additionally, we propose a multi-round prune-refine pipeline that can be applied to any pretrained 3D-GS model without changing its training pipeline. After pruning $\mathit{90\%}$ of Gaussians, a substantially higher percentage than previous methods, our PUP 3D-GS pipeline increases average rendering speed by $\mathit{3.56\times}$ while retaining more salient foreground information and achieving higher image quality metrics than existing techniques on scenes from the Mip-NeRF 360, Tanks & Temples, and Deep Blending datasets.
Figure 1: We prune the 3D Gaussian Splatting (3D-GS) reconstruction of the Deep Blending playroom scene from $2.65$M Gaussians to $0.265$M Gaussians using our PUP 3D-GS pipeline, accelerating rendering speed from $76.65$ FPS to $318.06$ FPS – a $\boldsymbol{4.15\times}$ speed-up on this scene– while preserving fine details. In comparison, LightGaussian, a recent high-performing post-hoc pruning pipeline for pretrained 3D-GS models, loses substantially more fine details than PUP 3D-GS and achieves a slower rendering speed of $261.27$ FPS.
@@ 1 Introduction
Novel view synthesis aims to render views from new viewpoints given a set of 2D images. Neural Radiance Fields (NeRFs) [20] use volume rendering to represent the 3D scene using a multilayer perceptron that can be used to render novel views. Although NeRF and its variants achieve high-quality reconstructions, they suffer from slow inference and require several seconds to render a single image. 3D Gaussian Splatting (3D-GS) [11] has recently emerged as a faster alternative to NeRF, achieving real-time rendering on modern GPUs and comparable image quality. It represents scenes using a large set of 3D Gaussians with independent location, shape, and appearance parameters. Since they typically consist of millions of Gaussians, 3D-GS scenes often have high storage and memory requirements that limit their viability on devices with limited resources.
Several recent works propose techniques that reduce the storage requirements of 3D-GS models, including heuristics for pruning Gaussians that have an insignificant contribution to the rendered images [4, 6, 16, 23, 14]. In this work, we propose a more mathematically principled approach for deciding which Gaussians to prune from a 3D-GS scene. We introduce a computationally feasible sensitivity score that is derived from the Hessian of the reconstructed error on the training images in a converged scene. Following the methodology in LightGaussian [4], we prune the scene using our sensitivity score and then fine-tune the remaining Gaussians. Our experiments demonstrate that this process can remove $80\%$ of the Gaussians in the base 3D-GS scene while preserving image quality metrics.
Moreover, our approach enables a second consecutive round of pruning and fine-tuning. As shown in Figure 1, our multi-round pruning approach can remove $90\%$ of the Gaussians in the base 3D-GS scene while surpassing previous heuristics-based methods like LightGaussian in both rendering speed and image quality. Our work is orthogonal to several other methods that modify the training framework of 3D-GS or apply quantization-based techniques to reduce its disk storage. We show that our PUP 3D-GS pipeline can be used in conjunction with these techniques to further improve performance and compress the model.
@@
    A post-hoc pruning technique, PUP 3D-GS, that can be applied to any pretrained 3D-GS model without changing its training pipeline.
@@
    A multi-round prune-refine approach that produces higher fidelity reconstructions than an equivalent single round of prune-refining.
@@ 2.1 Uncertainty Estimation
Several works estimate uncertainty in NeRF-based approaches that arise from sources like transient objects, camera model differences, and lightning changes [10, 25, 27, 18]. Other works focus on uncertainty caused by occlusion or sparsity of training views [19]. These approaches rely on an ensemble of models [31] or variational inference and KL Divergence [29, 30], requiring intricate changes to the training pipeline to model uncertainty.
BayesRays [7] uses Fisher information for post-hoc uncertainty quantification in NeRF-based approaches. Since NeRFs [19] represent the scene as a continuous 3D function, they compute the Hessian over a hypothetical perturbation field to estimate uncertainty. In contrast, our approach focuses on 3D Gaussian Splats [11] instead, and we directly compute the Fisher information using gradient information without relying on a hypothetical perturbation field.
@@ 2.2 Pruning Gaussian Splat Models
A recent pruning-based method, Compact-3DGS [14], proposes a learnable masking strategy to prune small, transparent Gaussians during training. EAGLES [6], which we apply our pipeline to in Appendix A.5, prunes Gaussians based on the least total transmittance per Gaussian. They also begin with low-resolution images, progressively increasing image resolution to reduce Gaussian densification during training, then quantize several attributes to reduce disk storage. LightGaussian [4] computes a global significance score for each Gaussian with heuristics, uses that score to prune the least significant Gaussians, and finally uses quantization to further reduce storage requirements. Mini-Splatting [5] and RadSplat [24] are concurrent with our work and introduce heuristics-based pruning components. However, they primarily focus on orthogonal components, such as alternative densification strategies and pretrained NeRF priors, for 3D-GS instead of pruning.
3 Background: 3D Gaussian Splatting
3D Gaussian Splatting (3D-GS) is a point-based novel view synthesis technique that uses 3D Gaussians to model the scene. The attributes of the Gaussians are optimized over input training views given by a set of camera poses $\mathcal{P}_{gt}=\{\phi_{i}\in\mathbb{R}^{3\times 4}\}_{i=1}^{K}$ and corresponding ground truth training images $\mathcal{I}_{gt}=\{\boldsymbol{I}_{i}\in\mathbb{R}^{H\times W}\}_{i=1}^{K}$. A sparse point cloud of the scene generated by Structure from Motion (SfM) over the training views is used to initialize the Gaussians. At fixed intervals during training, a Gaussian densification step is applied to increase the number of Gaussians in areas of the model where small-scale geometry is insufficiently reconstructed.
Each 3D Gaussian $\mathcal{G}_{i}$ is independently parameterized by position $x_{i}\in\mathbb{R}^{3}$, scaling $s_{i}\in\mathbb{R}^{3}$, rotation $r_{i}\in\mathbb{R}^{4}$, base color $c_{i}\in\mathbb{R}^{3}$, view-dependent spherical harmonics $h_{i}\in\mathbb{R}^{15\times 3}$, and opacity $\alpha_{i}\in\mathbb{R}$. From these, we define the set of all Gaussian parameters as:
@@ 3 Background: 3D Gaussian Splatting
For a given camera pose $\phi$, a differentiable rasterizer renders 2D image $I_{\mathcal{G}}(\phi)$ by projecting all Gaussians observed from $\phi$ onto the image plane. The color of each pixel $p$ in $I_{\mathcal{G}}(\phi)$ is given by the blending of the $\mathcal{N}$ ordered Gaussians that overlap it:
$C(p)=\sum_{i\in\mathcal{N}}\tilde{c}_{i}\tilde{\alpha}_{i}(p)\prod_{j=1}^{i-1}(1-\tilde{\alpha}_{j}(p)),$ (3)
where $\tilde{c}_{i}$ represents the view-dependent color calculated from the camera pose $\phi$ and the optimizable per-Gaussian color $c_{i}$ and spherical harmonics $h_{i}$, and $\tilde{\alpha_{i}}(p)$ represents the projected Gaussian value at $p$ times the Gaussian’s opacity $\alpha_{i}$.
@@
where the first term is a L1 residual loss and the second term is a SSIM loss.
@@ 4 Method
We define uncertainty in 3D-GS as the amount that a Gaussian’s parameters, such as its location and scale, can be perturbed without affecting the reconstruction loss over the input views. Concretely, this is the sensitivity of the error over the input views to that particular Gaussian. Given a loss function $L:\mathbb{R}^{\mathcal{G}}\to\mathbb{R}$ that takes the set of Gaussians $\mathcal{G}$ as inputs and outputs an error value over the set of training views, this sensitivity can be captured by the Hessian $\nabla^{2}_{\mathcal{G}}L$.
In the following subsections, we demonstrate how to compute this sensitivity and use it to decide which Gaussians to prune from the model. Directly computing the full Hessian matrix is intractable due to memory constraints. We demonstrate how to obtain a close estimate of the Hessian via a Fisher approximation in Section 4.1. We also find that only a block-wise approximation of the Hessian parameters is needed to quantify Gaussian sensitivity in Section 4.2, and that computing this over image patches is sufficient for pruning in Section 4.3. Finally, we find that multiple rounds of pruning and fine-tuning improves our performance over one-shot pruning and fine-tuning, giving our full PUP 3D-GS pipeline in Section 4.4.
4.1 Fisher Information Matrix
@@
On a converged 3D-GS model, the $||I_{\mathcal{G}}-I_{gt}||_{1}$ residual term of Equation 4 approaches zero, causing the second order term $(I_{\mathcal{G}}(\phi)-I_{gt})\nabla_{\mathcal{G}}^{2}I_{\mathcal{G}}$ in Equation 6 to vanish:
$\nabla_{\mathcal{G}}^{2}L_{2}=\sum_{\phi\in\mathcal{P}_{gt}}\nabla_{\mathcal{G}}I_{\mathcal{G}}(\phi)\nabla_{\mathcal{G}}I_{\mathcal{G}}(\phi)^{T}.$ (7)
This is identified as the Fisher Information matrix in related literature [7, 9]. We provide an explicit derivation of the Fisher Information matrix from this $L_{2}$ loss along with a Bayesian interpretation of our method in Appendix A.1. Note that $\nabla_{\mathcal{G}}I_{\mathcal{G}}$ is the gradient over only the reconstructed images, so our approximation only depends on the input poses $\mathcal{P}_{gt}$ and not the input images $\mathcal{I}_{gt}$.
@@ 4.2 Sensitivity Pruning Score
The full Hessian over the model parameters $\nabla_{\mathcal{G}}^{2}L\in\mathbb{R}^{\mathcal{G}\times\mathcal{G}}$ is quadratically large. However, we find that using only a subset of these parameters is sufficient for an effective sensitivity pruning score. For brevity, we remove the $\mathcal{P}_{gt}$ terms.
@@
$\boldsymbol{H}_{i}=\nabla_{\mathcal{G}_{i}}I_{\mathcal{G}}\nabla_{\mathcal{G}_{i}}I_{\mathcal{G}}^{T},$ (8)
where $\nabla_{\mathcal{G}_{i}}$ is the gradient with respect to the parameters of Gaussian $\mathcal{G}_{i}$. Intuitively, each block Hessian $\boldsymbol{H}_{i}$ measures the isolated impact that perturbing only $\mathcal{G}_{i}$’s parameters would have on the reconstruction error. To turn $\boldsymbol{H}_{i}$ into a sensitivity score ${\tilde{U}}_{i}\in\mathbb{R}$, we take its log determinant:
${\tilde{U}}_{i}=\log|\boldsymbol{H}_{i}|=\log|\nabla_{\mathcal{G}_{i}}I_{\mathcal{G}}\nabla_{\mathcal{G}_{i}}I_{\mathcal{G}}^{T}|.$ (9)
This captures the relative impact of all the parameters of Gaussian $\mathcal{G}_{i}$ on the reconstruction error. Specifically, it is a relative volume measure of the basin around the Gaussian parameters $\mathcal{G}_{i}$ in the second-order approximation of the function describing their impact on the reconstruction error.
We find that we can restrict the per Gaussian parameters even further and consider only the spatial mean $x_{i}$ and scaling $s_{i}$ parameters for an effective sensitivity score. Our final sensitivity pruning score $U_{i}\in\mathbb{R}$ is:
$U_{i}=\log|\nabla_{x_{i},s_{i}}I_{\mathcal{G}}\nabla_{x_{i},s_{i}}I_{\mathcal{G}}^{T}|.$ (10)
We speculate that only the mean and scaling parameters are needed because they capture the projective geometric invariances that exist between the 3D scene and the input views as shown in Figure 2. Rotations – the remaining geometric parameters – are excluded because they do not induce a change of 3D geometry when invariances are present. We ablate the choice of parameters in Section 6.3.